# Testing Noise Assumptions of Learning Algorithms

### v1

@@ C.1 Completeness
Since we have already proven the property of monotonicity under datapoint removal, we can assume without loss of generality that $M=U$. If not, the set $S$ can be obtained by first taking $U$ samples from $\mathcal{N}(0,I_{d})$ and then removing the last $U-M$ of them. If the $\mathcal{T}_{\text{spectral}}$ accepted the dataset before removing these points, then it will also accept it after these datapoints are removed.
@@ C.2 Soundness
[…]
@@ C.3 Bounding the $L_{2}$ sandwiching degree of the disagreement region
To prove Proposition C.4, we follow an exactly analogous approach as the one for Proposition B.5. We will need the following result from [KSV24b]:
Fact C.6 ([KSV24b]).
@@
Indeed, if $\frac{\log^{10}k}{\sqrt{k}}\leq\frac{1024}{t^{2}}$ then the corollary follows from Fact C.6. So all we need to do is to consider the other case. We see that either $t>1$ or $t<-1$ (since $k\geq 4)$. if $t>1$ we take $p_{\text{down}}(\mathbf{x})=0$ and $p_{\text{up}}(\mathbf{x})=\left(\frac{x}{t}\right)^{2}$. If $t<-1$, we take take $p_{\text{up}}(\mathbf{x})=1$ and $p_{\text{down}}(\mathbf{x})=1-\left(\frac{x}{t}\right)^{2}$. In either case, the polynomials $p_{\text{down}}$, $p_{\text{up}}$ form a pair of $(O(\min(\frac{\log^{10}k}{\sqrt{k}},\frac{1}{t^{2}})),1)$-sandwiching polynomials of degree $2$. ∎
@@
$\psi_{\text{up}}(\mathbf{x})=\begin{cases}1&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\geq k_{1}\epsilon$}\text{ or }\mathbf{v}\cdot\mathbf{x}=0,\\ \mathbbm{1}_{\mathbf{v}_{\perp}\cdot\mathbf{x}\tan\theta\leq-j\epsilon}&\text{\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\neq 0$ and }}\mathbf{v}\cdot\mathbf{x}\in\left[j\epsilon,(j+1)\epsilon\right)\text{ for }0\leq j\leq k_{1}-1,\\ 0&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$<0$}\end{cases}$
$\psi_{\text{down}}(\mathbf{x})=\begin{cases}0&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\geq k_{1}\epsilon$}\text{ or }\mathbf{v}\cdot\mathbf{x}=0\\ \mathbbm{1}_{\mathbf{v}_{\perp}\cdot\mathbf{x}\tan\theta<-(j+1)\epsilon}&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$\neq 0$ and }\mathbf{v}\cdot\mathbf{x}\in\left[j\epsilon,(j+1)\epsilon\right)\text{ for }0\leq j\leq k_{1}-1\\ 0&\text{if }\text{$\mathbf{v}\cdot\mathbf{x}$$<0$}\end{cases}$
@@ C.4 Miscellaneous Claims
Thus, we take the $\binom{d+1}{k_{2}}\times\binom{d+1}{k_{2}}$-matrix $W^{a,b}$ to have entries pairs of monomials $m_{1}$ and $m_{2}$ over $\mathbb{R}^{d}$ of degree at most $k$ and equal to additive $\beta^{\prime}$-approximations to ${\operatorname*{\mathbb{E}}}_{\mathbf{x}\sim\mathcal{N}}[m_{1}(\mathbf{x})m_{2}(\mathbf{x})\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}].$ Thus, the difference between $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}}[(\mathbf{x}^{\otimes k_{2}})(\mathbf{x}^{\otimes k_{2}})^{\top}\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}]$ and $W^{a,b}$ is a $\binom{d+1}{k_{2}}\times\binom{d+1}{k_{2}}$-matrix whose entries are bounded by $\beta^{\prime}$ in absolute value. Thus, the Frobenius norm of this difference matrix is at most $(d+1)^{k_{2}/2}\beta^{\prime}$, and therefore all eigenvalues of the matrix $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}}[m_{1}(\mathbf{x})m_{2}(\mathbf{x})\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}]-W^{a,b}$ are in $[-(d+1)^{k_{2}/2}\beta^{\prime},(d+1)^{k_{2}/2}\beta^{\prime}]$. Taking $\beta^{\prime}=\beta(d+1)^{-k_{2}/2}$, we can conclude that Equation C.17 holds.
@@ D.2 Completeness
We now argue that for a sufficiently large absolute constant $C$, the algorithm $\mathcal{\mathcal{A_{\text{Massart}}}}$ satisfies the completeness condition. In this subsection we assume that ${\bar{S}}$ is generated by ${\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}}$ where $f(\mathbf{x})=\operatorname{\mathsf{sign}}({\mathbf{v}^{*}}\cdot\mathbf{x})$ is an origin-centered halfspace and $\eta_{0}\leq 1/3$. We remind the reader that, for some function $\eta:\mathbb{R}^{d}\rightarrow[0,\eta_{0}]$, every time ${\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}}$ is invoked it generates an i.i.d. pair $(\mathbf{x},y)\in\mathbb{R}^{d}\times\left\{\pm 1\right\}$ where $\mathbf{x}$ is drawn from $\mathcal{N}(0,I_{d})$ and $y=f(\mathbf{x})$ with probability $\eta(\mathbf{x})$ and $-f(\mathbf{x})$ with probability $1-\eta(\mathbf{x})$. We would like to show that $\mathcal{A_{\text{Massart}}}$ accepts with probability at least $1-O(\delta)$.
@@
Additionally, recall that in this subsection we are assuming that ${\bar{S}}$ is generated by ${\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}}$. This implies that the elements of $S$ are generated i.i.d. from $\mathcal{N}(0,I_{d})$. Since the decision wheather each datapoint $\mathbf{x}$ in $S$is included into $S_{\text{augmented}}$ is made with probability $\eta_{0}$ independently from the actual value of $\mathbf{x}$, this implies the element of $S_{\text{augmented}}$ are distributed i.i.d. as $\mathcal{N}(0,I_{d})$ even conditioned on any specific value of $|S_{\text{augmented}}|$. This finishes the proof of the claim. ∎
@@
Recall that in step (1) of $A_{\text{Massart}}$ we used the algorithm of [DKTZ20a] (see 2.8) which implies that with probability at least $1-\delta$ we have $\measuredangle(\mathbf{v},{\mathbf{v}^{*}})\leq\frac{\epsilon^{3/2}}{10\sqrt{d-1}}$.

### v2

@@ C.1 Completeness
Since we have already proven the property of monotonicity under datapoint removal, we can assume without loss of generality that $M=U$. If not, the set $S$ can be obtained by first taking $U$ samples from $\mathcal{N}(0,I_{d})$ and then removing the last UMU-M of them. If the $\mathcal{T}_{\text{spectral}}$ accepted the dataset before removing these points, then it will also accept it after these datapoints are removed.
@@ C.2 Soundness
We say that a function $f:\mathbb{R}^{d}\rightarrow\left\{0,1\right\}$ has $(\epsilon,B)$-sandwiching degree of at most $k$ in $L_{2}$ norm under $\mathcal{N}(0,I_{d})$ with respect to a partition $\mathcal{C}$ of $\mathbb{R}^{d}$ if the function $f$ is $\epsilon$-sandwiched in $L_{2}$ norm under $\mathcal{N}(0,I_{d})$ between $\sum_{A\in\mathcal{C}}\left(p_{\text{down}}^{A}\mathbbm{1}_{A}\right)$ and $\sum_{A\in\mathcal{C}}\left(p_{\text{up}}^{A}\mathbbm{1}_{A}\right)$, where $p_{\text{up}}^{A}$ and $p_{\text{down}}^{A}$ are degreek-k polynomials over $\mathbb{R}^{d}$whose coefficients are bounded by $B$ in absolute value.
@@ C.3 Bounding the $L_{2}$ sandwiching degree of the disagreement region
To prove Proposition C.4, we follow an exactly analogous approach as the one for Proposition B.5. We will need the following result from [klivans2023testable]:
Fact C.6 ([klivans2023testable]).
@@
Indeed, if $\frac{\log^{10}k}{\sqrt{k}}\leq\frac{1024}{t^{2}}$ then the corollary follows from Fact C.6. So all we need to do is to consider the other case. We see that either $t>1$ or t<1t<-1 (since $k\geq 4)$. if $t>1$ we take $p_{\text{down}}(\mathbf{x})=0$ and $p_{\text{up}}(\mathbf{x})=\left(\frac{x}{t}\right)^{2}$. If t<1t<-1, we take take $p_{\text{up}}(\mathbf{x})=1$ and $p_{\text{down}}(\mathbf{x})=1-\left(\frac{x}{t}\right)^{2}$. In either case, the polynomials $p_{\text{down}}$, $p_{\text{up}}$ form a pair of $(O(\min(\frac{\log^{10}k}{\sqrt{k}},\frac{1}{t^{2}})),1)$-sandwiching polynomials of degree $2$. ∎
@@
[…]
@@ C.4 Miscellaneous Claims
Thus, we take the $\binom{d+1}{k_{2}}\times\binom{d+1}{k_{2}}-$matrix $W^{a,b}$ to have entries pairs of monomials $m_{1}$ and $m_{2}$ over $\mathbb{R}^{d}$ of degree at most $k$ and equal to additive $\beta^{\prime}$-approximations to ${\operatorname*{\mathbb{E}}}_{\mathbf{x}\sim\mathcal{N}}[m_{1}(\mathbf{x})m_{2}(\mathbf{x})\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}].$ Thus, the difference between $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}}[(\mathbf{x}^{\otimes k_{2}})(\mathbf{x}^{\otimes k_{2}})^{\top}\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}]$ and $W^{a,b}$ is a $\binom{d+1}{k_{2}}\times\binom{d+1}{k_{2}}-$matrix whose entries are bounded by $\beta^{\prime}$ in absolute value. Thus, the Frobenius norm of this difference matrix is at most $(d+1)^{k_{2}/2}\beta^{\prime}$, and therefore all eigenvalues of the matrix $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{N}}[m_{1}(\mathbf{x})m_{2}(\mathbf{x})\cdot\mathbbm{1}_{a\leq\mathbf{x}\cdot\mathbf{v}<b}]-W^{a,b}$ are in $(d+1)^{k_{2}/2}\beta^{\prime},(d+1)^{k_{2}/2}\beta^{\prime}]$. Taking $\beta^{\prime}=\beta(d+1)^{-k_{2}/2}$, we can conclude that Equation C.17 holds.
@@ D.2 Completeness
We now argue that for a sufficiently large absolute constant $C$, the algorithm $\mathcal{\mathcal{A_{\text{Massart}}}}$ satisfies the completeness condition. In this subsection we assume that $\bar{S}$} is generated by $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}$} where $f(\mathbf{x})=\operatorname{\mathsf{sign}}({\mathbf{v}^{*}}\cdot\mathbf{x})$ is an origin-centered halfspace and $\eta_{0}\leq 1/3$. We remind the reader that, for some function $\eta:\mathbb{R}^{d}\rightarrow[0,\eta_{0}]$, every time $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}$} is invoked it generates an i.i.d. pair $(\mathbf{x},y)\in\mathbb{R}^{d}\times\left\{\pm 1\right\}$ where $\mathbf{x}$ is drawn from $\mathcal{N}(0,I_{d})$ and $y=f(\mathbf{x})$ with probability $\eta(\mathbf{x})$ and $f(\mathbf{x})$ with probability $1-\eta(\mathbf{x})$. We would like to show that $\mathcal{A_{\text{Massart}}}$ accepts with probability at least $1-O(\delta)$.
@@
Additionally, recall that in this subsection we are assuming that $\bar{S}$} is generated by $\mathsf{EX}^{\mathsf{Massart}}_{\mathcal{N},f,\eta_{0}}$}. This implies that the elements of $S$ are generated i.i.d. from $\mathcal{N}(0,I_{d})$. Since the decision wheather each datapoint $\mathbf{x}$ in SSis included into $S_{\text{augmented}}$ is made with probability $\eta_{0}$ independently from the actual value of $\mathbf{x}$, this implies the element of $S_{\text{augmented}}$ are distributed i.i.d. as $\mathcal{N}(0,I_{d})$ even conditioned on any specific value of $|S_{\text{augmented}}|$. This finishes the proof of the claim. ∎
@@
Recall that in step (1) of $A_{\text{Massart}}$ we used the algorithm of [diakonikolas2020learning] (see 2.8) which implies that with probability at least $1-\delta$ we have $\measuredangle(\mathbf{v},{\mathbf{v}^{*}})\leq\frac{\epsilon^{3/2}}{10\sqrt{d-1}}$.